OpenStudy (anonymous):

Segment EF is the midsegment of trapezoid ABCD. Find the length of segment AB if segment EF = 120 and segment DC is 100. Enter the numerical answer only. co031-1.jpg

OpenStudy (goformit100):

O.o^

jimthompson5910 (jim_thompson5910):

hint: EF = (AB+DC)/2

OpenStudy (anonymous):

co031-1.jpg wasn't supposed to be there that's just the picture :P

OpenStudy (anonymous):

so the answer is 110??

jimthompson5910 (jim_thompson5910):

EF = (AB+DC)/2 120 = (x+100)/2 120*2 = x+100 x = ???

OpenStudy (anonymous):

hahah im pretty lost

jimthompson5910 (jim_thompson5910):

120 times 2 = what number

OpenStudy (anonymous):

...240

jimthompson5910 (jim_thompson5910):

240 = x+100 x = ???

OpenStudy (anonymous):

lol 140

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

wow ;p so my answer is 140?!

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yay thanks

jimthompson5910 (jim_thompson5910):

you're welcome